OpenStudy (anonymous):

What is the standard form equation of the line passing through (5,-3) with a slope of -3

OpenStudy (campbell_st):

the line is y +3 = -3(x - 5) expanding gives y + 3 = -3x + 25 standard form is ax + by + c = 0 just rewrite the above equation

OpenStudy (anonymous):

another way is y=mx+b -3=5*-3+b sovling for b you get b=12 so -3= 5*-3+12 y=x*-3+12 y-x*-3=12

OpenStudy (anonymous):

Thanks Tim that is one of the answers thanks alot

OpenStudy (anonymous):

@campbell_st, @timo86m I strongly feel you should refrain from giving complete answers. I think the formula is more sufficient than enough \[y - y_1 = {(y_2 - y_1) \over (x_2-x_1)} (x-x_1)\] where (x_1, y_1)-->point 1 on the line (x_2, y_2)-->point 2 on the line Now he should have solved it. I am saying this based on Openstudy Guidelines. Make sure you have a check at them

OpenStudy (anonymous):

IDK if we gave complete answers. she can mess around with the equation if she want to make it equal 0 or have a constant on rhs

OpenStudy (anonymous):

Well, then in that case @timo86m , you never ever made mistake while learning for the first time.??Our duty must be to teach and not give complete answers

OpenStudy (anonymous):

well depending on the person. I like to learn by having the answer then work on my own to reach that answer :)

OpenStudy (anonymous):

Ok . Let @Directrix have his say

OpenStudy (anonymous):

depends on attitude of person. If they want the answer for easy grade and never learning anything then thhat is bad. But to have the answer as a guide to check if your work is right then that is good.

Directrix (directrix):

@Vanessa173 From above in thread: >standard form is ax + by + c = 0 ^^ is incorrect You may want to recheck your answer options in light of the following: Standard Form for Linear Equation Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

@timo86m , agree to your point of view but @Vanessa173 was only interested in the final answer as I can see from her statements So basically, we are bunch of people who are answering (probably) giving her answer out for the test :P

OpenStudy (anonymous):

@Directrix , agreed :D

OpenStudy (anonymous):

lol very rarely are there take home test. Probably only on online classes but i bet the are timed and waiting here for an answer would suck.

OpenStudy (anonymous):

LOL, idk . as far as I know we shouldn't be giving answers to the people here as the moderators will serve you with a warning here. Have a look at Open study guide lines. I have also got warned once. So that's why alerting you.